OpenStudy (anonymous):

integral of e^(-x^2)x dx

OpenStudy (johnweldon1993):

\[\large \int e^{-x^2}xdx\] If we let \(\large u = -x^2\) that means \(\large du = -2xdx\) Now notie we need to get \(\large xdx\) but now we have \(\large -2xdx\) so lets divide everything by -2 \[\large -\frac{1}{2} \int e^{u}du\] And we know the integral of e^x = e^x so \[\large -\frac{1}{2}e^u +c\] Now lets just replace back in 'u' in terms of 'x' \[\large -\frac{1}{2}e^{-x^2} + c\]